### v1

# Continual Learning in Medical Imaging Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

Abstract
Medical imaging analysis has witnessed remarkable advancements even surpassing human-level performance in recent years, driven by the rapid development of advanced deep-learning algorithms. However, when the inference dataset slightly differs from what the model has seen during one-time training, the model performance is greatly compromised. The situation requires restarting the training process using both the old and the new data which is computationally costly, does not align with the human learning process, and imposes storage constraints and privacy concerns. Alternatively, continual learning has emerged as a crucial approach for developing unified and sustainable deep models to deal with new classes, tasks, and the drifting nature of data in non-stationary environments for various application areas. Continual learning techniques enable models to adapt and accumulate knowledge over time, which is essential for maintaining performance on evolving datasets and novel tasks. Owing to its popularity and promising performance, it is an active and emerging research topic in the medical field and hence demands a survey and taxonomy to clarify the current research landscape of continual learning in medical imaging analysis. This systematic review paper provides a comprehensive overview of the state-of-the-art in continual learning techniques applied to medical imaging analysis. We present an extensive survey of existing research, covering topics including catastrophic forgetting, data drifts, stability, and plasticity requirements. Further, an in-depth discussion of key components of a continual learning framework such as continual learning scenarios, techniques, evaluation schemes, and metrics is provided. Continual learning techniques encompass various categories, including rehearsal, regularization, architectural, and hybrid strategies. We assess the popularity and applicability of continual learning categories in various medical sub-fields like radiology and histopathology. Our exploration considers unique challenges in the medical domain, including costly data annotation, temporal drift, and the crucial need for benchmarking datasets to ensure consistent model evaluation. The paper also addresses current challenges and looks ahead to potential future research directions.
Keywords Continual Learning  $\cdot$ Drift  $\cdot$ Domain Shift  $\cdot$ Concept Drift  $\cdot$ Medical Image Analysis  $\cdot$ Histopathology  $\cdot$ Radiology
1 Introduction
In the evolving field of medical imaging analysis, the dynamic nature of healthcare data poses a critical challenge for the generalizability of the machine learning/deep learning models to new data/domains. The data-driven approaches have challenges due to the limited availability and accessibility of sufficiently large and diverse medical data for training. Additionally, the source variability, scanner manufacturer, imaging protocol, and diverse population groups make the medical data heterogeneous, this adds a bias and discrepancies between training and testing datasets, thus leading to performance degradation. In order to handle the model generalizability issues, domain adaptation methods become popular and aim to transfer knowledge from one domain to other unseen data sources or domains. However, domain adaptation poses unique challenges due to the sensitive and complex nature of healthcare data Azad et al. (2022). The most common associated issues are the limited availability of labeled medical data for training, heterogeneity in data sources contributing to significant domain shifts, clinical disparities and population variances, and inter-rater variabilities. Also, biases present in the source domain data can propagate to the target domain, so ensuring fair and unbiased predictions across diverse patient populations becomes a critical concern. Moreover, medical data encompasses various modalities, including imaging, electronic health records, and genomic data, making the task of adapting models to handle multimodal data and ensuring interoperability exceptionally complex.
Continual learning (CL) - continuous adaptation to new information, has emerged as an important dimension in enhancing the performance and reliability of medical image analysis systems. Continual learning models are employed for predictive analytics, especially in situations where clinical outcomes can be automatically obtained and incorporated into the algorithm. This capability enhances the model’s predictive power by learning from real-time patient data. CL methods can also be employed to utilize the multi-modality dataset for a better interpretability and analysis. In recent years, an increasing number of continual learning methods have been explored and proposed in various subareas of computer vision tasks. Figure 1 presents the paper distributions based on different aspects of CL and shows the growth in exploration in this domain through the rising number of publications over the years. In this paper, we discuss the various aspects of continual learning, particularly considering its application and implications in the medical domain. Here, we aim to contribute to the ongoing discourse on adapting and improving machine learning models for sustained effectiveness in the dynamic healthcare landscape. By emphasizing continual learning, we recommend models that not only demonstrate robust performance initially but also possess the ability to evolve and improve over time. Continual learning in medical image analysis represents a powerful approach to developing intelligent systems that can evolve, learn, and adapt to the complexities of healthcare, ultimately contributing to improved patient outcomes and enhanced clinical decision-making. We discuss the limitations and challenges of the existing methods and explore the methods/techniques that can be utilized for developing a robust algorithm. A thorough search of the existing body of literature reveals the groundbreaking nature of our work as the first comprehensive survey of continual learning techniques applied to medical imaging analysis. This scholarly endeavor aims not only to contribute novel insights but also to establish a foundational reference for researchers, offering a roadmap that can guide future exploration and incite scholarly interest in the academic community. The primary contributions of our academic pursuit are elucidated below:
$\bullet$ Pioneering in its scope, this survey paper provides the first comprehensive exploration of continual learning applications in the field of medical imaging. Our focus extends to delivering a thorough overview encompassing all pertinent papers and elucidating details regarding well-known methods in medical image analysis.
$\bullet$ We introduce a rigorous categorization of continual learning models within the academic community, presenting a systematic taxonomy that categorizes research based on different continual learning strategies. Our classification discerns between various continual learning techniques, such as rehearsal, regularization, architectural, and hybrid methods. Additionally, we contextualize these techniques within various medical sub-fields, offering a nuanced academic perspective.
$\bullet$ Beyond application-centric discussions, our exploration delves into the scholarly challenges and open issues surrounding continual learning in medical imaging analysis. By addressing academic intricacies, including data annotation costs, temporal drift, and the necessity for benchmarking datasets, we contribute to the scholarly discourse. Additionally, we identify emerging academic trends that give rise to open questions, shaping the trajectory of future academic research in continual learning applied to medical imaging analysis.
Motivation and Uniqueness of this Survey
Over the past few decades, continual learning approaches in computer vision tasks have seen substantial progress, leading to numerous survey papers exploring deep continual learning models for computer vision tasks Qu et al. (2021); Wang et al. (2023a); De Lange and Tuytelaars (2021); Mai et al. (2022). While some of these surveys focus on specific applications, such as classifications De Lange and Tuytelaars (2021), others take a more general approach to concepts and practical perspectives Wang et al. (2023a). Notably, none of these surveys specifically addresses the applications of continual learning techniques in medical imaging, leaving a significant gap in the literature. We believe that insights from successful continual learning models in vision can be beneficial for the medical community, guiding the retrospective analysis of past and future research directions in continual learning Verma et al. (2023); Lee and Lee (2020). Continual learning has proven its potential in developing unified and sustainable deep models capable of handling new classes, tasks, and the evolving nature of data in non-stationary environments across various application areas. Our survey aims to bridge the gap by providing valuable insights that can assist medical researchers, including radiologists, in adopting up-to-date methodologies in their fields. In our survey, we analyze various sources of drifts in medical data, defining continual learning scenarios in medical imaging. We present a multi-perspective view of continual learning by categorizing techniques into rehearsal, regularization, architectural, and hybrid strategies-based methods. The discussion extends beyond applications, encompassing underlying working principles, challenges, and the imaging modality of the proposed methods. We emphasize how this additional information can aid researchers in consolidating literature across the spectrum. A concise overview of our paper is illustrated in Figure 6.
Search strategy
To conduct a thorough literature search, we follow the same strategy presented in Azad et al. (2023a, b) and utilized DBLP, Google Scholar, and Arxiv Sanity Preserver, employing custom search queries to retrieve scholarly publications related to our topic-continual learning. Our search query was (continual learning | medical | sequence of tasks) (segmentation | classification | medical | lifelong learning) . These platforms allowed us to filter results into categories such as peer-reviewed journal papers, conference or workshop proceedings, non-peer-reviewed papers, and preprints. We carefully filtered our search results through a two-step process: initially screening based on the title and abstract, followed by a subsequent screening of the full text. This approach was instrumental in eliminating false positives, ensuring that only relevant content was included in our study.
Paper organization
All the abbreviations used in this manuscript along with their expansion are tabulated in Table 1. A background about various sources of drift in medical data, the catastrophic forgetting issue, continual learning, and its application is provided in Section 2. Then a thorough description of various kinds of continual learning scenarios found in the medical domain is elaborated in Section 3. Further, we cover various categories of continual learning techniques and their applicability in the medical domain via Section 4. Section 5 provides details about the level of supervision required in different kinds of proposed frameworks. Section 6 provides comprehensive practical information such as the experimental setups, training process, and evaluation metrics for measuring the plasticity and stability of continual learning frameworks. Section 7 discusses the current challenges in the continual learning literature and future directions. Eventually, the last section provides the conclusion of the survey.
Table 1: Abbreviations and their expansion
Acronym Full name
CL Continual Learning
MAS Memory Aware Synapses
VQA Visual Question-Answering
sMRI structural Magnetic Resonance Imaging
H&E Hematoxylin and Eosin
CMR Cardiac Magnetic Resonance
WBC White Blood Cells
CIS Class Incremental Scenario
DIS Domain Incremental Scenario
TIS Task Incremental Scenario
EWC Elastic Weight Consolidation
DWC Distributed Weight Consolidation
SurfDist Symmetric surface Distance
Acronym Full name
IoU Intersection over Union
AUC Area Under ROC Curve
MAE Mean Absolute Error
CMR Cardiovascular Magnetic Resonance
AP Average Precision
DSC Dice Similarity Coefficient
HD95 95% Hausdorff Distance
ASD Average symmetric Surface Distance
FROC Free-response Receiver-Operating Characteristic
ROC Receiver-Operating Characteristic
MSD Mean Surface Distance
ASSD Average Symmetric Surface Distance
MCR Mean Class Recall
2 Background
2.1 Medical data drifts in non-stationary environment
In clinical practices, the data distribution evolves over time, reflecting the dynamic nature of the healthcare environment. Inconsistencies in data collection procedures across different healthcare settings or institutions contribute to data drift. Also, the introduction of new medical imaging devices, diagnostic tools, and data acquisition techniques leads to a shift in the technological landscape. Moreover, the advancement in medical research and the discovery of new diseases/treatment methods raise the understanding of healthcare. This new knowledge can influence the characteristics of medical data that cause shifts in the underlying distribution.
The medical data drifts can have significant implications for the performance and reliability of machine learning models. Traditional machine learning often relies heavily on static data and feature engineering, where human experts manually select the relevant features. In the case of data drift, these handcrafted features may become less informative, and the model may struggle to adapt to new patterns. Understanding the sources of drift, such as inconsistencies in data collection and technological advancements, is crucial for developing robust models. Proactive strategies, including regular model updates and continuous monitoring, are essential to ensure that machine learning models remain effective and reliable in navigating the evolving healthcare landscape. By addressing medical data drift and developing machine learning models that can adapt to the ever-changing nature of clinical data, ultimately enhancing patient care and outcomes.
2.2 Catastrophic forgetting
Throughout a lifetime, a human brain continuously acquires knowledge, and learning new concepts or tasks has no detrimental effect on previously learned ones. Instead, learning several closely related concepts even boosts the learning of all associated ones. In contrast, artificial neural networks, although inspired by the human brain, often suffer from ‘catastrophic forgetting’, a tendency to overwrite or forget the knowledge acquired in the past upon learning a new concept McCloskey and Cohen (1989); Ratcliff (1990). This can be attributed to the fact that the model entirely optimizes for the given dataset. In other words, a model with optimized weights for a task A, when trained on a new task B, will freely optimiz the existing weights to meet the objectives in task B, which may now no longer be optimal for the previous task A. This can be a significant challenge, especially in scenarios where an AI system is expected to learn and adapt to a stream of tasks or datasets over time. Catastrophic forgetting in neural networks is an interesting phenomenon that has attracted lots of attention in recent research Goodfellow et al. (2013).
2.3 Continual learning overview
A naive solution to deal with catastrophic forgetting can be retraining the model on old and new data from scratch, each time new data with drifted distribution or classes are encountered. This process mostly gives the desired classification or segmentation performance however causes an intense burden on computing and storage requirements and hence impractical for deployment. Additionally, the retraining process requires storing the past data and thus causes privacy violations which can be a major bottleneck of such a strategy in health-care applications.
Alternately, Continual Learning (CL), also termed as ‘continuously learning’ or ‘incremental learning’ or ‘lifelong learning’ has emerged as a promising solution in various fields to deal with the catastrophic forgetting issue in neural networks. It helps in efficiently leveraging existing knowledge and incorporating new information without the need for extensive retraining. The primary goal of CL is to develop techniques and strategies that allow a neural network to learn new tasks while retaining knowledge of previous tasks. In other words, it aims to enable the network to continually adapt to new information without completely erasing or degrading its performance on earlier tasks. Overall, CL helps to address the issue of catastrophic forgetting and minimizes the need for additional resources to store historical data. CL offers a range of strategies and methods, such as regularization (constraining weight update to avoid forgetting learnt concepts), rehearsal (partially using old data to replay with current data), and architectural modifications (reserving or partitioning network for different tasks), to help neural networks remember and consolidate knowledge from past tasks. These strategies help prevent or reduce catastrophic forgetting and improve the generalization ability of the model.
Figure 1: A coarse level flowchart for designing a CL pipeline
2.4 Continual learning pipeline
The design of a CL pipeline is illustrated in  Figure 1 which provides an overview of the key components and stages involved in the construction of the CL pipeline. Given a problem statement, first, we need to identify which CL scenarios it falls under, i.e., whether there is a possibility of domain shifts in future data, inclusion of new classes, or the end application, i.e., the task itself may change. For example, if we want to develop a breast cancer classification model to be able to work on the H&E dataset from different centers then the datasets across centers may have drift and hence fall into the domain incremental scenario of CL. Detailed information about CL scenarios is provided in Section 3. Once we have identified the CL scenario, training and testing datasets need to be prepared to mimic a continual stream of datasets arriving one after another. The sequence of datasets is frequently referred to as tasks or episodes. Since the term ‘task’ is also used for end application and thus may be challenging to get the true interpretation of the term, hence, we use the ‘episode’ term to refer to the sequence of datasets in a CL pipeline. For each episode, separate training and testing data is prepared, thus to develop and evaluate a CL model for four sequential datasets, the pipeline requires four train-test pairs. Once the datasets are ready, a CL strategy suitable to the task at hand is identified and deployed in any off-the-shelf deep classification or segmentation model. There are various CL strategies, some offering privacy-preserved learning, while some offer better performance but at the cost of more resources, storage, and privacy violations as they store some past data. Typically, the model is trained on the first training episode and evaluated on testing data from all the episodes. After this the training shifts to the next episode where the inclusion of partial training data from the previous episode is possible. Here, after training, the evaluation is again done on all the testing data and the process repeats until the last episode. Application-specific performance metrics (e.g., accuracy, dice similarity coefficient, etc.) on testing data of each episode are observed and analyzed over the sequence. Then, we can compute various metrics on top of it to quantify forgetting and forward transfer. Lastly, the CL framework is evaluated against state-of-the-art (SOTA), algorithms offering upper and lower bounds of performance. Joint or cumulative training gives the highest average performance whereas naively finetuning on the current episode gives the lowest performance .
Figure 2: Ratio of CL-based research for downstream applications
Figure 3: CL-based research contributions over the year
2.5 Application of continual learning
Continual learning, also known as incremental learning, has numerous applications across various domains due to its ability to mitigate catastrophic forgetting. The efficiency and robustness of CL in real-time scenarios derive from its ability to adapt, reduce computational overhead, and address the challenges of dynamic data. This makes it a valuable approach for applications where time constraints, adaptability, and efficiency are of great importance. In the medical domain, CL has been widely explored for various segmentation and classification applications (Figure 2) and continuously exhibited its merits over static models as reflected by the increasing number of research contributions (Figure 3). CL can improve the diagnosis and decision-making ability in a resource-constrained environment. Further, it is beneficial in real-time monitoring of patients, telemedicine, and maintaining a dynamic knowledge base. This section presents the most prominent applications of continual learning in medical settings in terms of the number of publications.
Radiology and imaging: Radiological imaging techniques are continually advancing, requiring models to adapt to new technologies and methodologies. For CT, research has investigated the impact of scanners and reconstruction parameters on both machine learning predictions and human annotations. The findings indicate that the variability introduced by different scanners has a detrimental effect on radiomics  Mackin et al. (2015)  Prayer et al. (2021) and other imaging features. Therefore, it is imperative to take into account this scanner variability when developing machine learning models. Continual learning facilitates the dynamic adjustment of models to evolving imaging techniques, ensuring accurate interpretations  Li and Hoiem (2018)  Zhang et al. (2023a). Table 13 and Table 14 present the selected works for continual learning based classification and segmentation in the medical domain.
Disease progression modeling: Continual learning is crucial in tracking the progression of diseases such as Alzheimer’s Disease (AD). Models need to adapt to new patient data over time, incorporating the latest diagnostic criteria and treatment strategies Li et al. (2020). One research proposed a method for AD progression in a continual learning manner that respects the longitudinal data sets coming in sequence and ensures equal prediction accuracy for future visits  Zhang and Wang (2019).
Drug discovery: Continual learning can accelerate drug discovery by continuously integrating newly discovered chemical compounds, pharmacological data, and clinical trial results to predict drug interactions and efficacy. A recent work presented a continual learning-based model called Multi-Scale Temporal Convolutional Networks-based AntiBacterial Peptide prediction (MSTCN-ABPpred). This model is designed for the classification and discovery of antibacterial peptides (ABPs). While MSTCN-ABPpred can dynamically adapt and retrain based on predicted ABPs and non-ABPs within protein sequences, it does have limitations in terms of providing information about the identified ABPs’ targets, haemotoxicity, cytotoxicity, and minimum inhibitory concentration (MIC) against various bacteria  Singh et al. (2023).
Figure 4: Ratio of CL-based works for different incremental scenarios
3 Continual learning scenarios
Depending upon what kind of change in the sequence of data is expected, the existing CL scenarios can be broadly categorized into five categories, viz., (a) data-incremental, (b) class-incremental, (c) task-incremental, (d) domain-incremental, and (e) hybrid. Figure 4 shows the percentages of the above-mentioned incremental CL scenarios exploited in recent years. The statistics reflect that out of four major settings, domain-incremental is highly explored with a 51.9% ratio followed by class-incremental settings having 26.9% works and data-incremental being the least challenging and hence rarely explored for different medical image analysis applications.
Table 2: List of various data incremental scenarios in literature
Reference (year) Application (# Episodes) Description
Ravishankar et al. (2019)
Pneumothorax identification from X-ray images (2 Ep.)
4 batches each of 2K were created from a subset of ChestXRay (Wang et al. 2017)
Kaustaban et al. (2022)
Histopathology data-based tumor classification (4 Ep.)
Ep. was created in data incremental fashion from colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes, (i.e., all 9 classes are present in each episode)
3.1 Data incremental scenarios
Data incremental scenario involves the continuous learning process where the model keeps adapting to incoming data streams that come from the same source. This scenario doesn’t involve dealing with entirely new categories or significant changes in data patterns, making it generally the least challenging among all other continual learning scenarios. In the medical domain, especially digital pathology, a common practice is for expert pathologists to annotate a dataset in multiple stages or batches Kaustaban et al. (2022). Each of these annotation batches can be thought of as representing samples from the same underlying data distribution, ensuring some consistency in the data source across different batches. Thus, data incremental scenarios deal with the condition where the samples of a dataset are sequentially made available. Owning to its simplicity, this category is not very popular. Only a few classification applications are explored in this scenario, as mentioned in Table 2. For example, in the breast cancer classification application, Chen and Tang (2022) set up data incremental scenarios on the BreakHis Spanhol et al. (2015) dataset. More and more samples from each of the malignant and benign classes are added over the subsequent episodes.  Ravishankar et al. (2019) try to exhibit a scenario for pneumothorax detection in a hospital where data arrives in an incremental fashion. They sequentially fed 2000 samples in 4 subsequent episodes from an X-ray dataset for Pneumothorax identification consisting of a total of 8000 samples.
Table 3: List of various class incremental scenarios in literature
Reference (year) Application (# Episodes) Description
Ozdemir et al. (2018)
MRI based Humerus and scapula segmentation (2 Ep.)
Incrementally adding anatomical structure, private data
Ozdemir and Goksel (2019)
MRI based knee segmentation (2 Ep.)
SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
Li et al. (2020)
Dermoscopic images based skin disease classification (4-20 Ep.)
(a) CIS for Skin8 dataset having total 8 classes on which 2 classes per episode used, (b) CIS for Skin40 dataset from ISIC2019 (Tschandl et al. 2018) consists of 40 classes of skin disease images collected from the internet, here 2/5/10 classes per episode used in 3 exp, (c) non-medical dataset CIFAR100 with 100 classes, here 5/10/20 classes per episode used in 3 exp
Liu et al. (2022)
Class incremental segmentation of abdomen organs (4 Ep.)
4 Ep. created for incremental segmentation of liver, spleen, pancreas, right kidney, and left kidney organs from CT datasets (Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
Derakhshani et al. (2022)
CIS: disease classification (4 Ep.)
4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (taken from MedMNIST repository (Yang et al. 2023a))
Akundi and Sivaswamy (2022)
Chest X-ray classification in CIS (5 Ep.)
5 classes from CheXpert dataset
Kaustaban et al. (2022)
Histopathology data based tumor classification (4 Ep.)
4 Ep. were created in CIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total 9 classes
Chen and Tang (2022)
Histopathology data-based breast cancer classification (4 Ep.)
4 Ep. created in CIS from BreakHis dataset
Zhang et al. (2023b)
(a) CT based abdomen, gastrointestinal, and other organ segmentation (7, 3, 4 classes in 3 Ep.), (b) abdomen segmentation to liver tumor segmentation task (13 classes, 1 class in 2 Ep.)
(a) Abdomen segmentation from JHH (Xia et al. 2022) (private) (b) 13 class abdomen segmentation dataset as BTCV (Landman et al. 2015) and liver tumor segmentation data as LiTS (Bilic et al. 2023)
Ji et al. (2023)
3D CT scan based whole-body organs segmentation (4 Ep. in 2 permutations of episodes)
TotalSegmentator (Wasserthal et al. 2023) (103 classes: various organs), 3 private datasets: ChestOrgan (31 classes: chest scans), HNOrgan (13 classes: head and neck scans), EsoOrgan (1 class: esophageal cancer)
Chee et al. (2023)
Cancer classification (3,5 Ep.), diabetic retinopathy classification (2 Ep.), skin lesions classification (3,4 Ep.)
(a) Histopathology data based human colorectal cancer classification: CCH5000 (Kather et al. 2016) (total 8 classes, Exp=(4,1,1,1,1), Exp=(4,2,2)), (b) diabetic retinopathy (DR) classification using retinal images: EyePACS (kag ) (total 5 classes: no DR, mild DR, moderate DR, severe DR and proliferative DR, Exp=(3,2)), (c) pigmented skin lesions classification: HAM10000 (Tschandl et al. 2018) (total 7 classes: Exp=(4,1,1,1), Exp=(4,2,1))
Zhang et al. (2023c)
Disease classification in CIS (4-10 Ep.)
(a) skin lesions classification with Skin8 dataset (International Skin Imaging Collaboration (ISIC) (Tschandl et al. 2018)) having 8 classes distributed over 4 Ep., (b) Path16, a pathology image collection from various public sources: total 16 classes distributed over 7 episodes, (c) CIFAR100: non-medical images, 100 classes distributed over 5-10 episodes
(Bai et al. 2023)
Surgical visual-question localized-answering
EndoVis18, EndoVis17, M2CAI
Sadafi et al. (2023)
Microscopic images based WBC classification (4 Ep.)
With 3 different datasets, 3 separate class-incremental scenarios each having 4 episodes: (a) (3,3,3,4) classes from Matek-19 (Matek et al. 2019), (b) (3,3,3,4) classes from INT-20, and (c) (3,2,2,3) classes from Acevedo-20 (Acevedo et al. 2020)
3.2 Class incremental scenarios
Class Incremental Scenario (CIS) refers to the situation where we aim to adapt the model to accommodate novel classes from new data streams. Given that each batch of these unseen classes offers knowledge that largely differs from the previous batches, this scenario is anticipated to be the most challenging among continual learning scenarios. An instance of this challenge is incorporating new types of tissues into a tumor classification model or introducing novel cell types into a model designed for AI-assisted diagnosis in cell detection. In this scenario, there is only one incremental task that gets new additions of classes in subsequent episodes. Some assume mutually exclusive classes whereas some overlapping classes.
Mainly multi-class datasets with more than 5-6 classes are considered to create a class incremental scenario. There can be an equal or variable number of classes across the episodes. Chen and Tang (2022) propose a breast cancer classification model where they iteratively include malignant and benign classes with one sub-class from each category. This scenario is very popular for segmentation applications where the model may get demand to segment more and more classes over time. An exhaustive list of class incremental scenarios explored in literature is provided via Table 3. References along with detailed segmentation or classification applications are tabulated.
Table 4: List of various task incremental scenarios in literature
Reference (year) Application (# Episodes) Description
Baweja et al. (2018)
MRI based normal brain structures segmentation (2 Ep.)
[…]
Ravishankar et al. (2019)
Chest x-ray view classification (2 Ep.)
Shift from task (4ch vs. PLAX) to task (2ch vs. PSAX)
Kaustaban et al. (2022)
Histopathology data based tumor classification (4 Ep.)
4 Ep. were created in TIS from Colorectal cancer dataset CRC (Kather et al. 2019) having a total of 9 classes
Derakhshani et al. (2022)
Disease classification (4 Ep.)
4 Ep. created from each of the datasets in TissueMNIST, OrganaMNIST, PathMNIST, BloodMNIST (MedMNIST repository (Yang et al. 2023a))
Bera et al. (2023)
Three tasks involving different organs for three different binary segmentation tasks (3 Ep.)
MRI data for binary prostate, spleen, hippocampus segmentation: Promise12 (Litjens et al. 2014) (prostate) $\rightarrow$ MSD (NCI ) (spleen) $\rightarrow$ Drayd (Denovellis et al. 2021) (hippocampus)
3.3 Task incremental scenarios
Task Incremental Scenario (TIS) comes into the picture when we have a multi-task problem and a single adaptive model is desired. Each task is considered an episode. Thus each episode has disjoint label space. However, there is ambiguity in the literature with task incremental scenarios as the other scenarios (class and domain) are frequently referred to as task incremental scenarios by different research communities. For example, Ravishankar et al. (2019) create the first episode having two chest x-ray views (2ch vs. PLAX) and then the next episode with two other views (4ch vs. PSAX) as a task incremental scenario which can also be a class incremental scenario. Further another example in this line can be found in the work by Baweja et al. (2018). Authors create a task incremental scenario with two episodes for brain MRI segmentation application where the $1^{st}$ episode was multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the $2nd$ episode was the segmentation of white matter lesions. If we have multiple tasks that are very different then it would certainly be the task incremental; however, if the tasks are more close then it is merely a design choice whether to treat them as class incremental or task incremental. Another work Kaustaban et al. (2022) explores domain incremental scenario as organ shift for tumor classification problem where the first episode is a colon cancer dataset with 9 classes (CRC dataset) and the next episode is a breast cancer dataset with 2 classes (PatchCam dataset). Similar is the case with Sadafi et al. (2023) where WBC classification from three datasets, each having a different number of classes is regarded as domain incremental which can also be regarded as task-incremental. Derakhshani et al. (2022) curate class incremental and task incremental scenarios from the same dataset taken from MedMNIST Yang et al. (2023a) where a dataset is divided into non-overlapping classes leading to 4 episodes. If the model is evaluated only on the learnt class in the specific episode then they regard it as task incremental else if evaluated on cumulative classes from all seen episodes then class incremental scenario.
Kaustaban et al. (2022) say that any of the scenarios within data, class, and domain incremental can be treated as task incremental, provided that each incoming data stream (episode) is treated as a distinct task. For a comprehensive overview, we have tabulated all the works along with their claimed task incremental scenario in Table 4.
Further, it is essential to consistently provide prior information about which specific task (referred to as task identity, i.e., task ID) the test data pertains to, and predictions are made accordingly. However, if task IDs are not provided then this category is termed as a task-free scenario which is a more challenging one.
Table 5: List of various domain shift scenarios in literature
Reference (year) Application (# Episodes) Description
McClure et al. (2018)
Cross-site (4 Ep.)
sMRI based Axial and sagittal brain segmentation: datasets from 4 sites: HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017) in different order
Karani et al. (2018)
cross-scanners, hospitals, or acquisition protocols (4 Ep.)
MR brain segmentation on various combination using 5 datasets from different source: Human Connectome Project (Van Essen et al. 2013), Alzheimer’s Disease Neuroimaging Initiative (ADNT ), Autism Brain Imaging Data Exchange (Di Martino et al. 2014) and Information eXtraction from Images (IXI )
van Garderen et al. (2019)
Low vs. high-grade (2 Ep.)
MR data based glioma segmentation: 2018 BraTS Challenge (low and high-grade), an in-house dataset with non-enhancing low-grade
Venkataramani et al. (2019)
varying in disease type, intensity patterns (2) & contrast
X-ray lung segmentation: Montgomery (Jaeger et al. 2014) as source, JSRT (Shiraishi et al. 2000) and Pneumoconiosis (private) as two target domains
Ravishankar et al. (2019)
Inter-subject variability (2 Ep.)
(a) Chest x-ray view classification (4ch vs. PLAX) from adult to pediatric as two domains (b) chest x-ray view classification (2ch vs. PSAX) from adult to pediatric as two domains
Lenga et al. (2020)
Cross-sites (2 Ep.)
Chest x-ray classification on datasets from 2 source: ChestX-ray14 dataset from NIH Clinical Center and MIMIC-CXR dataset from the Beth Israel Deaconess Medical Center
Özgün et al. (2020)
Different age ranges, MRI field strengths & presence of pathologies or motion artifacts (total 4 Ep.)
Brain MRI segmentation datasets: 4 Ep. were created from 3 datasets: CANDI (Kennedy et al. 2012) (1 Ep.), ADNI (Jack Jr et al. 2008) (1 Ep.), and MALC (Asman and Landman 2013) (2 Ep. based on age)
Hofmanninger et al. (2020a)
Total 3 Ep. (a) different scanner parmeter (b) target-shift
Chest CT synthetic classification task: (a) first to second episode have change in scanner parameter (Hofmanninger et al. 2020b) (b) target shift is introduced intentionally in the third episode by imprinting a synthetic target structure in the form of a cat on random locations, rotations and varying scale in 50% of the samples.
Morgado et al. (2021)
Appearance and body-parts (2 Ep.)
Dermatological imaging modality classification (full-body, anatomic, macroscopic, and dermoscopic classes): For each of the class author considered disjoint images in the two episodes, for example, in the second episode, legs and arms images were considered for full-body class, images containing hands or feet only considered for anatomic, images containing face region for macroscopic class, and pink colored images for dermoscopic class.
Srivastava et al. (2021)
Cross-sites (3 Ep.)
Chest x-ray classification: NIH Chest-X-rays14, PadChest, and CheXpert
Memmel et al. (2021)
Multi-scanner (2 Ep.)
MRI-based Hippocampal Segmentation: 3 datasets as 2018 Medical Segmentation Decathlon challenge (Simpson et al. ), Scientific Data (Kulaga-Yoskovitz et al. 2015), and Alzheimer’s Disease Neuroimaging Initiative (Boccardi et al. 2015): 2 datasets together in 1st episode and then 3rd dataset in 2nd episode, different such combinations were experimented
Perkonigg et al. (2021)
Multi-scanners (3 Ep.)
Brain age estimation: IXI(IXI ) (Philips Gyroscan Intera 1.5T, Philips Intera 3.0T scanner), OASIS3 (LaMontagne et al. 2019) (Siemens TrioTim 3.0T scanner)
Bayasi et al. (2021)
Multi-site (6 Ep.)
Skin lesion image classification: HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
Zhang et al. (2021)
Cross-sites (4-6 Ep.)
(a) Prostate segmentation: MRI datasets across 6 sites: RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (b) optic cup and disc segmentation: public multi-site fundus image datasets from 4 sources (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
Perkonigg et al. (2022)
Multi-scanner (4 Ep.)
(a) CMR based Cardiac segmentation (4 Ep.): Cardiac (Campello et al. 2021), (b) CT based Lung nodule detection (4 Ep.): LIDC (Setio et al. 2017), + LNDb challenge (Pedrosa et al. 2019), (c) MRI based Brain Age Estimation segmentation (4 Ep.): IXI(IXI ) + OASIS-3 (LaMontagne et al. 2019)
Derakhshani et al. (2022)
Across organs or modality (4 Ep.)
disease classification: 4 epi using 4 dataset ordered as: BloodMNIST, OrganaMNIST, PathMNIST, TissueMNIST (MedMNIST repository)
Karthik et al. (2022)
Cross-center (8 Ep.)
Sclerosis lesions segmentation from brain MRI datasets described in Kerbrat et al. (Kerbrat et al. 2020)
González et al. (2022)
(a) Cross-domain (5 Ep.) (b) manual image contrast change (5 Ep.)
MRI hippocampus segmentation: (a) HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019) (b) MRI hippocampus segmentation: data transformation applied using TorchIO library on Decathlon (Simpson et al. 2019) as intensity rescaling, affine transformations, rotation, translation
Ranem et al. (2022)
Cross-site (3 Ep.)
MRI based binary hippocampus segmentation: Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
Kaustaban et al. (2022)
(a) Multi-organ (2 Ep.) (b) scanning protocol (5 Ep.)
(a) Tumor classification: CRC (a colon cancer dataset with 9 classes), PatchCam (a breast cancer dataset with 2 classes) (b) 5 domain scenario were simulated by changing H&E composition in CRC dataset
Shu et al. (2022)
(a) Imaging protocol (2 Ep.) (b) multi-source (3 epi)
(a) Low and high-quality fundus images (retinal images) from EyeQ dataset as domain shift condition (Fu et al. 2019), (b) fundus disease classification (AMD, DR, glaucoma, myopia, and normal classes): ODIR (ODIR ), R&R (RIADD (RIADD ) + REFUGE (REFUGE )), and iSee (Fang et al. 2020)
Li et al. (2022)
Multi-scanner vendors (4 Ep.)
CMR based Cardiac segmentation of left ventricle, right ventricle, and left ventricle myocardium: 4 scanner vendors (Siemens, Philips, General Electric, Cannon) from M&Ms (Campello et al. 2021) dataset
Sadafi et al. (2023)
Cross-sites (3 Epi)
WBC classification datasets: Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Bera et al. (2023)
(a) Cross-center (4 Ep.) (b) cross-center (2 Ep.)
(a) Binary prostate segmentation: Prostate158 (Adams et al. 2022) $\rightarrow$ NCI-ISBI (NCI ) $\rightarrow$ Promise12 (Litjens et al. 2014) $\rightarrow$ Decathlon (Antonelli et al. 2022) (b) binary hippocampus segmentation: Drayd (Denovellis et al. 2021) $\rightarrow$ HarP (Boccardi et al. 2015))
Zhu et al. (2023)
(a) Cross-site (6 Ep.) (b) cross-site & cross-modality (2 Ep.) (c) same-site & cross-modality (2 Ep.)
(a) Binary prostate segmentation from T2-weighted MRI scans collected from 6 sites (12–30 scans/site) (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015) (b) multi-class (liver, left and right kidneys, and spleen) abdominal segmentation between 30 CT and 20 MRI T2-SPIR scans (c) muscle segmentation of 13 lower-leg muscles and bones between 30 MRI T1 and 30 mDixon scans
Bándi et al. (2023)
Cross-organ (3 Ep.)
Histopathology data based cancer detection from breast (CAMELYON16, CAMELYON17), colon (private), and head-neck (private) datasets
3.4 Domain incremental scenarios
Domain Incremental Scenario (DIS) is the most popular and frequently observed category of CL scenarios for medical applications. This scenario aligns with the idea of continual learning in a changing environment where the datasets from different domains (e.g., research sites, hospitals, imaging modality, image acquisition protocol, etc.) are incrementally encountered over time which thus involve covariance shift-induced discrepancies in data.
In clinical applications, this might encompass changes in factors like the methods used for tissue processing and staining, the demographics of the patient population, or the types of scanning instruments employed, among other variables. Thus, continuous modifications in diagnostic techniques result in alterations in the appearance of medical images. Factors such as the brand of the scanning machine, the method used to create the images, radiation dosage, and specific settings in the scanning process, including the use of contrast agents, can all influence how the images look. These changes in image characteristics, which occur independently of the actual biological content being scanned lead to domain shifts.
Even if we have a single task and a fixed number of classes, these shifts can pose a challenge for the static deep model used in clinical settings because these shifts can quickly make existing models outdated and less effective. Then a CL model aims to continuously update itself by incorporating new data streams that come from distributions that have shifted over time.
There have been plenty of attempts to develop CL models and evaluate different kinds of domain incremental scenarios. All the domain incremental scenarios are not equi-hard; some works use a simpler level of domain shift whereas some with severe domain shift. To provide a comparative view, we tabulate all the domain incremental scenarios considered in literature through Table 5.
Table 6: List of various hybrid CL scenarios in literature
Reference (year) Application (# Episodes) Description
Liu et al. (2023a)
Brain tumor segmentation (3 Ep.)
CIS with domain shift conditions: incrementally learn tumor core (BraTS2013 (Menze et al. 2014) dataset), the enhancing tumor (TCIA (Clark et al. 2013) dataset), and edema (CBICA (Bakas et al. 2018) dataset) structures via three datasets, each following different data distributions
Sadafi et al. (2023)
WBC classification (3 Ep.)
DIS (multi-site) + CIS using 3 WBC classification datasets having different number of classes: CIS on Matek-19 (Matek et al. 2019) then shift to INT-20 dataset and use it in CIS manner then shift to Acevedo-20 (Acevedo et al. 2020) dataset and use in CIS manner
Yang et al. (2023b)
MedMNIST (Yang et al. 2023a) based disease classification (4 Ep.)
CIS+DIS: 3 datasets (PathMNIST, DermaMNIST, and OrganAMNIST) as source domain and then 3 more datasets (RetinaMNIST, BreastMNIST, BloodMNIST) as domain shift but classes are added in class incremental fashion with 1 class at a time
Wang et al. (2023b)
CIS (a) segmentation for endoscopy (2-3 Ep.) (b) surgical instrument segmentation (2 Ep.)
Incrementally adding segmentation structure from EDD2020 (Ali et al. 2021, 2020) dataset which includes 5 classes as Barrett’s esophagus, cancer, high-grade dysplasia, polyp, and suspicious. Initially, 3-4 classes are in 1st episode, and then 2 or 1 classes are incrementally added in subsequent episodes. (b) EndoVis18 (Allan et al. 2020) and EndoVis17 (Allan et al. 2019) datasets as two episodes having some overlapping classes
Liu et al. (2023a)
MRI-based brain tumor segmentation (4 Ep.)
CIS+DIS Incrementally adding more structures to segmentation task from dataset acquired from different vendors/centers
3.5 Simulated or Hybrid Settings
It may not be always the case that the model will always observe a particular kind of incremental scenario, for example, domain incremental condition. The CL model developed for handling domain incremental conditions may not perform well for a class incremental condition. Thus, there are attempts to evaluate the same model on different CL scenarios; however, a separate evaluation is followed for each kind of incremental CL scenario. In contrast, having a set of episodes with a mix of CL scenarios is close to real-life conditions. Thus it becomes important to also test the performance in hybrid settings. An attempt in direction is made by Sadafi et al. (2023) where other than pure domain incremental and class incremental, a hybrid incremental scenario is also considered. The authors create a sequence of episodes with class and domain-incremental cases on multi-site WBC classification datasets. They create a long sequence of episodes with a total of 12 episodes. First, they set up a class incremental scenario within the Matek-19 dataset by adding 3-4 classes in each subsequent episode up to a total of four episodes. After this INT-20 dataset was considered in the same fashion, contributing 4 more episodes. After this, a third domain i.e., a new dataset Acevedo-20 was introduced in 4 episodes in class incremental fashion.
Another novel direction for having a mix of scenarios is considered by Liu et al. (2023a). In contrast to traditional class incremental settings where class-incremental methods are developed without considering data shift, Liu et al. (2023a) incrementally add more structures for brain tumor segmentation application (class incremental) where the incremental structure can be from distinct sites/scanners/MRI modality (domain shift conditions).
Further, there are some attempts to curate shifts in datasets and various incremental scenarios instead of collecting actual datasets reflecting the situation. For example, Kaustaban et al. (2022) simulated domain shift in H&E stain due to distinct appearance caused by different staining protocols. This dataset is designed to simulate real-world scenarios, where data shifts occur due to differences in scanners, stainers, reagents, and other factors. They considered 9 classes from CRC Kather et al. (2019) which is a H&E stain-based colorectal cancer classification dataset. Each class is divided into 5 disjoint sets for 5 episodes. The first set, regarded as Domain1 is without any alteration and the next 4 sets for each class undergo various changes to create domain shift. A short description of 4 domain shifts is as follows: Domain 2 (increased stain intensity, simulating, for example, concentration increase of the eosin and/or hematoxylin solutions, each with a different extent of change). Domain 3 (decreased eosin stain intensity, simulating, for example, slides prepared from many years ago with fading stain). Domain 4 (change of hue, simulating, for example, change of reagent manufacturer, scanner or stainer) Domain 5 (change of hue and saturation, simulating, for example, change of reagent manufacturer, scanner or stainer). finally, authors used this augmented dataset for setting class-incremental, domain-incremental, data-incremental, as well as task-incremental scenarios.
Another line of research popularly uses MNIST or similar well-defined datasets instead of collecting real datasets for the evaluation of the CL model. In the medical field, there is a recently released (2021) MNIST-like collection of biomedical images offering dozens of datasets, termed MedMNIST Yang et al. (2023a). Derakhshani et al. (2022) consider 4 MedMNIST datasets including BloodMNIST, OrganaMNIST, PathMNIST, and TissueMNIST as 4 domains and curate three CL scenarios with it. These datasets are multi-class (8-11 classes in each) disease classification datasets from different imaging modalities and organs. Each dataset was split to have disjoint classes leading to a total of 4 episodes in each dataset. Task and class incremental scenarios were set up for each dataset separately. In the task incremental scenario, the aim was to evaluate only the classes from the current episode whereas in class incremental, the aim was to evaluate all seen classes. Thus the class incremental scenario is difficult than task incremental. They term the datasets generated from different sites/imaging protocols as cross-domain incremental scenario, instead of domain-incremental. Here the evaluation is done across the 4 datasets (BloodMNIST, OrganaMNIST, PathMNIST, and TissueMNIST), i.e., each dataset is treated as an episode. Further authors presented domain-agnostic and domain-aware settings in this. Domain-agnostic case is more difficult than domain-aware as the domain ID is not explicitly provided in domain-agnostic.
Figure 5: Popularity of different CL strategies for medical image analysis
Figure 6: The proposed CL taxonomy categorizes various CL models according to their underlying design principles. The abbreviations used are as follows: ER for Experience-reply, GR for Generative-reply, LR for Latent-reply, PF for Prior-focused, DF for Data-focused, and MBO for Memory Buffer Optimization. 1. Hofmanninger et al. (2020a), 2. Perkonigg et al. (2021), 3. Perkonigg et al. (2022), 4. Karthik et al. (2022), 5. Bera et al. (2023), 6. Li et al. (2022), 7. Srivastava et al. (2021), 8. Baweja et al. (2018), 9. McClure et al. (2018), 10. van Garderen et al. (2019), 11. Özgün et al. (2020), 12. Zhang et al. (2021), 13. Zhang et al. (2023d), 14. Chen and Tang (2022), 15. Zhu et al. (2023), 16. Liu et al. (2023a), 17. Karani et al. (2018), 18. González et al. (2022), 19. Memmel et al. (2021), 20. Bayasi et al. (2021), 21. Ozdemir et al. (2018), 22. Ozdemir and Goksel (2019), 23. Chee et al. (2023), 24. Li et al. (2020), 25. Ravishankar et al. (2019), 26. Liu et al. (2022), 27. Akundi and Sivaswamy (2022), 28. Shu et al. (2022), 29. Zhang et al. (2023b), 30. Sadafi et al. (2023), 31. Yang et al. (2023b),
4 Continual learning strategy
Owing to frequently encountered domain shifts in medical applications, CL strategies aim to tackle performance drop due to domain shifts and thus might become a necessity in clinical applications Pianykh et al. (2020); Lee and Lee (2020). All the available strategies for preventing catastrophic forgetting can be broadly categorized into three main categories: (a) rehearsal-based approaches where a small memory is used to store previous episode data in some form, (b) regularization-based methods aim to control weight update to minimize forgetting previous learning, (c) architectural-based methods mainly aim to keep some network parameters isolated for each episode, and (d) hybrid category which offers various combination of any of these three categories. Figure 6 presents a taxonomy of CL techniques and their adaptation in the medical field by various works. Further, the sub-technique category for each work is also mentioned in the figure. Further, a pie chart showing the ratio of works falling in the above-mentioned categories is shown via Figure 5. It indicates the popularity of using one or more strategies together in hybrid settings owing to better classification or segmentation performances in sequential learning. Now we discuss each of the categories in detail in the following sections.
Table 7: List of various rehearsal based works in literature
Reference (year) CL scenario & application CL technique
Hofmanninger et al. (2020a)
Chest CT classification: DIS (3 Epi.)
Experience replay with a dynamic memory size where samples are ranked by computing Gram matrix on last activation maps
Srivastava et al. (2021)
Chest x-ray classification: DIS(3 Epi.)
Latent replay
Perkonigg et al. (2021)
Brain age estimation: DIS (3 Epi.)
Experience replay
Perkonigg et al. (2022)
DIS: three separate experiments for 3 applications: cardiac segmentation (4 Epi.), lung nodule detection (4 Epi.), and brain age estimation (4 Epi.)
Experience replay
Karthik et al. (2022)
DIS on MRI data for brain sclerosis lesions segmentation (8 Epi.)
Random samples from each episode kept for experience replay
Li et al. (2022)
DIS: CMR based Cardiac segmentation across 4 scanners
Generative replay: style-oriented replay
Bera et al. (2023)
TIS (binary segmentation of prostate, spleen, and hippocampus), DIS (binary hippocampus segmentation across 2 datasets), DIS (binary prostate segmentation across 4 datasets)
Experience replay based rehearsal strategy: store max K exemplars (half of the exemplars based on fraction of positive class content inside the image and other half based on gradient variation)
Wang et al. (2023b)
(a) Segmentation for endoscopy: CIS (2-3 Epi.) (b) surgical instrument segmentation (2 Epi.)
Generative replay
4.1 Rehearsal based
This category of methods aims to approximate and recover old data distributions to augment with the new episode data. Typically there is a memory buffer to store data samples from past episodes which are then used for replaying during the learning of a new episode in order to mitigate catastrophic forgetting of previously learnt knowledge. Here the samples can be original images Rebuffi et al. (2017); Karthik et al. (2022) or deep features Van de Ven and Tolias (2019) or generated pseudo samples Shin et al. (2017) and can be selected via various heuristics and stored in memory buffer.
Based on the content of the memory buffer, the methods in this category can be broadly divided into the following sub-categories viz., (i) experience replay-based, (ii) generative replay-based, (iii) latent replay-based, and (iv) memory buffer optimization.
4.1.1 Experience replay-based
In experience replay, a few past exemplars (images) are stored in a small memory buffer. The main challenge here is to design a strategy to select important exemplars that will be stored in the limited storage to fully exploit the memory buffer.
For natural image classification, Rebuffi et al. (2017) proposed a popular experience-based method called incremental Classifier and Representation Learning (iCaRL) which is also highly explored for various medical applications, e.g., histopathology tumor classification Kaustaban et al. (2022), disease classification Derakhshani et al. (2022), etc. The iCaRL has a fixed memory buffer condition where a subset of the most representative examples is maintained for each class, aiming to carry the most representative information of the class in the learned feature space. The distance between data instances in the latent feature space is used to update the memory buffer. During representation learning, both the stored samples and current task samples are utilized for training. During inference, a nearest-mean-of-exemplars classification strategy is used to assign the label to the class with the most similar prototype. The original iCaRL method requires all data from the new task to be trained together. To address this limitation and enable the new instances from a single task to come at different time steps, Chaudhry et al. (2019) proposed Experience Replay (ER), which uses reservoir sampling to randomly sample a certain number of data instances from a data stream of unknown length, and store them in the memory buffer.
There has been further exploration to select the most representative samples for replay in medical applications. Bera et al. (2023) propose a simple sample selection technique for binary segmentation with a memory bank of $K$ samples, equi-distributed over the classes. Half of the exemplars are selected based on the occupancy of the positive class inside the image, i.e., the higher the content of the positive class, the higher the chance of being selected; rest other half are selected based on their contribution to the learning process, i.e., if the gradient variation is more for a sample then it is hard to learn and hence important. Hofmanninger et al. (2020a) propose to store samples based on their degree of uniqueness inferred using the Gram matrix computed on activations from the last convolution layer of the deep model. They evaluated the framework on chest CT data with synthetically generated domain shifts for classification application. Further, Perkonigg et al. (2021, 2022) propose to combine rehearsal-based continual learning and active learning which indicates important samples that need to be annotated rather than annotating all the samples. They detect domain shifts using a memory buffer for outliers, prompting the labeling of informative samples for model adaptation.
4.1.2 Latent replay-based
While experience replay requires storing the past samples in raw form, it can cause serious privacy violations in critical medical applications. Therefore a less concerning direction i.e., storing features instead of raw images is also explored. For continual chest x-ray classification application in domain shift conditions, Srivastava et al. (2021) explore leveraging vector-quantization to store and replay hidden representations under memory constraints.
4.1.3 Generative replay-based
This category of approaches emerged as an alternative to experience replay which highly violates privacy concerns with storing past samples and high memory buffers demands. Here instead of a memory buffer of actual samples, there is a generative model that can generate samples, latent representation, both, or other information related to past tasks. This category of methods is also related to continual learning of generative models where their incremental update is required. Hence, an additional requirement here is the continuous updating of the generative adversarial network. Since actual samples are not stored for replay, this category of approaches is also termed pseudo-rehearsal-based approaches.
To generate the past data, Li et al. (2022) employ a style-oriented replay module which includes a base generative model trained on the first arrived domain and a style bank to record style adjustments for successive domains. Then they incorporate the replayed past data to jointly optimize the model with current data to alleviate catastrophic forgetting. For continual semantic segmentation of MRI data emerging from different institutions, Memmel et al. (2021) train a model by using all the simultaneously available datasets. Their method disentangles content from domain information through adversarial training, resulting in domain-invariant content representations. Other methods follow sequential arrival of a dataset, i.e., one dataset at a time whereas this work requires at least two datasets to start the feature disentanglement learning.
4.1.4 Memory buffer optimization
Experience replay-based approaches might overfit the stored sub-samples and seem to be bounded by joint training. Alternatively, memory buffer optimization-based approaches following constrained optimization provide solutions that have more scope for both plasticity and stability. The core idea is to perform gradient projection. For example, the Gradient Episodic Memory (GEM) Lopez-Paz and Ranzato (2017) approach corrects the gradient computed on a mini-batch during stochastic gradient descent by utilizing an exemplar set. This prevents changes that could degrade the performance of the network on the exemplar set. Further, the A-GEM approach Chaudhry et al. (2018a) relaxes the projection in one direction computed by randomly selected samples from the memory buffer where both the buffer and sampling size are tunable hyper-parameters. There are some early attempts in this area in the medical realm. For example, for the fundus disease diagnosis problem, Shu et al. (2022) present a gradient regularisation approach for preventing forgetting and a replay-oriented consistency calculation method combined with a subspace weighting strategy to promote model adaptability. The authors employ gradient regularisation in conjunction with a replay-oriented technique. The replay-oriented technique adaptively improves the update for incremental domains without carefully choosing exemplars for replay, whereas gradient regularisation preserves information from previous ones.
Table 8: List of various regularization based works in literature
Reference (year) CL scenario & application CL technique
Baweja et al. (2018)
MRI based normal brain structures segmentation: TIS (2 Ep.)
EWC
McClure et al. (2018)
Axial and sagittal brain segmentation DIS (4 Ep.)
Distributed Weight Consolidation (DWC)
van Garderen et al. (2019)
MR data based glioma segmentation: DIS (2 epi)
EWC
Özgün et al. (2020)
Brain MRI segmentation: DIS (4 epi)
Learning rate regularization built over memory aware synapses technique
Chen and Tang (2022)
CIS for histopathology-based breast cancer classification (4 Epi)
Regularization: EWC
Zhang et al.; Zhang et al. (2023d; 2021)
DIS for (a) MRI based prostate segmentation, (b) DIS for retinal image based optic cup and disc segmentation
Shape and semantics-based selective regularization to penalize changes of parameters with high joint shape and semantics-based importance
Zhu et al. (2023)
(a) DIS: binary prostate segmentation from T2-weighted MRI scans collected from 6 sites, (b) cross-site and cross-modality 4-class abdominal segmentation between CT and MRI scans, (c) same-site cross-modality muscle segmentation of muscles and bones between MRI and mDixon scans
Knowledge distillation
Liu et al. (2023a)
Incremental brain tumor segmentation: CIS with domain shift (3 epi)
Pseudo label-based knowledge distillation
4.2 Regularization-based
Rehearsal methods are quite popular due to comparative better performance than other categories; however, the assumption for the availability of past data makes it less suitable for a medical application. In contrast, Regularization-based approaches Kirkpatrick et al. (2017); Schwarz et al. (2018); Zenke et al. (2017) avoid storing examples and mainly add a regularization term in the loss function or regularize the learning rate to penalize model updates that could lead to large deviation from an existing model, thus avoiding forgetting of learned knowledge.
Regularization-based methods can be categorized into data-focused and prior-focused approaches. Data-focused approaches Silver and Mercer (2002) distill the knowledge of old tasks to enhance the CL capabilities of the present model, whereas prior-based approaches such as Zenke et al. (2017); Kirkpatrick et al. (2017); Aljundi et al. (2018) define importance weights for the network’s parameters. Based on these weights, a regularization loss is introduced that penalizes the shift of important parameters.
4.2.1 Prior-focused regularization
Prior-focused methods estimate the importance of all neural network parameters, used as prior when learning from new data. During the training of subsequent tasks, larger changes to important parameters are penalized. Elastic Weight Consolidation (EWC) Kirkpatrick et al. (2017), initially designed for reinforcement learning in Atari games emerged as the first to establish the technique. EWC aims to constrain parameters of the model that are critical for performing previous tasks during the learning of the new tasks. It uses the Fisher information matrix to calculate parameter importance for a given domain. These parameters are then regularized to prevent catastrophic forgetting.
In the medical domain, Baweja et al. (2018) are the first to explore EWC Kirkpatrick et al. (2017) to address the issue of catastrophic forgetting in neural networks when sequentially learning two distinct segmentation tasks (normal brain structure and white matter lesion). Specifically, the first task consists of multi-class segmentation of cerebrospinal fluid, grey matter, and white matter, and the second task consists of segmentation of white matter lesions. The study demonstrates that EWC effectively reduces catastrophic forgetting in this challenging medical imaging context. Further, van Garderen et al. (2019) adopt EWC for glioma segmentation on different datasets in domain-shift arising from low and high-grade glioma in different datasets. In the context of histopathology breast cancer classification task, Chen and Tang (2022) showcase EWC capabilities for class incremental setting. McClure et al. (2018) introduce Distributed Weight Consolidation (DWC) as a continual learning method to consolidate weights of separate neural networks trained on independent datasets. Further, Zhang et al. (2023d) propose to compute the importance matrix jointly based on shape and semantics information in context for continual medical segmentation. The shape-based importance measures how sensitive a parameter is to shape properties in the images. and semantics-based importance measures a parameter’s sensitivity to reliable semantic predictions, ensuring that noisy or uncertain semantics do not influence the learning process. Then the selective regularization scheme is applied to penalize updates to model parameters with high joint importance weights. This ensures that critical shape and semantic knowledge from previous sites is not overwritten or forgotten.
Another prior-focused strategy is Synaptic Intelligence (SI) which alleviates catastrophic forgetting by allowing individual synapses (i.e., neurons) to estimate their importance for solving a learned task. Similarly to EWC, the approach penalizes changes to the most relevant synapses so that new tasks can be learned with minimal forgetting Zenke et al. (2017). A disadvantage is the need to distribute some extra parameters per weight in addition to their value, but in terms of data size, this is far less than providing the exemplars.
Apart from EWC and SI, various other methods have been contributed. Popularly, the Memory Aware Synapses (MAS) approach Aljundi et al. (2017) calculates the importance of weights with a model of Hebbian learning in the biological system, which relies on the sensitivity of the output function and can hence be utilized in an unsupervised manner. For medical segmentation, a MAS-inspired learning rate regularization approach was contributed by Özgün et al. (2020) for sequential training on different domains. It involves reducing the learning rate for important parameters to prevent forgetting, offering a more direct approach compared to the surrogate loss used in MAS.
4.2.2 Data-focused regularization
Data-focused regularization method Li and Hoiem (2017) aims to distill knowledge from a model trained on the previous tasks to the model trained on the new task in order to consolidate previously learned knowledge. Typically, the previous model acts as a teacher and the current model as a student while adopting knowledge-distillation Gou et al. (2021) to avoid catastrophic forgetting. Usually, all the old data should be available in knowledge distillation which is not the case with CL; therefore, as an alternative, few old data or the current data or the generated data are explored for the same. The loss function has an additional distillation loss for replayed data. Each input is replayed with a soft target obtained using the stored model.
Silver and Mercer (2002) first proposed to use previous task model outputs given new task input images, mainly for improving new task performance. Later, in natural image classification application, Li and Hoiem (2017) re-introduced the concept as LwF Li and Hoiem (2017) method. A copy of the previous model parameters is stored before learning the new task and then it is used to get the soft labels for the new task as the target for the classifiers of the previous tasks. The available ground truth is used as the target for the new task classifier. LwF has been adopted in various medical applications like incremental brain MRI segmentation Ozdemir et al. (2018), chest x-ray classification Lenga et al. (2020), disease classification Derakhshani et al. (2022), histopathology tumor classification Kaustaban et al. (2022), etc.
Other works in natural image domain Jung et al. (2016); Zhang et al. (2020) have been introduced with LwF-related ideas, however, it has been shown that this strategy is vulnerable to domain shift between tasks Aljundi et al. (2017). It was pointed out that LwF would not result in good performance if the data distributions between different tasks are quite diverse Rannen et al. (2017). To overcome this, Rannen et al. (2017) facilitate incremental integration of shallow auto-encoders to constrain task features in their corresponding learned low dimensional space. Hence, they further trained an autoencoder for each task to learn the most important features corresponding to the task and used it to preserve knowledge.
Table 9: List of various architecture based works in literature
Reference (year) CL scenario & application CL technique
Karani et al. (2018)
MR brain segmentation: DIS (4 Ep.)
Domain specific batch normalization layer
Bayasi et al. (2021)
Skin lesion image classification: DIS (6 Ep.)
Architecture based (fixed but partitioned network)
Memmel et al. (2021)
Brain MRI Hippocampal Segmentation: DIS (2 Ep.)
Architecture (feature disentanglement through adversarial training)
González et al. (2022)
DIS with (2 Ep.) for MRI based Hippocampus Segmentation
Architecture based approach based on out of distribution detection concept: maintain multivariate Gaussians for all past learning created on batch normalization layer
4.3 Architecture-based
Architecture-based methods, also termed parameter isolation-based methods typically assign different parameters in a network to each task. This can be achieved by either fixing the architecture or dynamically extending the network Mallya and Lazebnik (2018); Hung et al. (2019); Mallya and Lazebnik (2018); Fernando et al. (2017); Yoon et al. (2017). Fixed architectures are limited by the network’s capacity, whereas dynamic architectures need more memory with every new task. Most of these methods use different parts of the network for each task which requires task identity during inference, but this identity information is usually unavailable.
Some works Yan et al. (2021a) aim to dynamically extend the feature extractor module by learning task-specific branches in it. Rusu et al. (2016) propose to use a dynamic architecture that blocks any changes to the network trained on previous knowledge and expands the architecture by allocating sub-networks with a fixed capacity to be trained with the new data and thus it keeps a pool of pre-trained models, one for each learned task. In contrast, Aljundi et al. (2017) propose a network of experts where each expert is a model trained given a specific task and a set of gating autoencoders that learn a representation for the task at hand, and, at inference time, automatically forward the test sample to the relevant expert. Another dynamic expandable model direction given by Yoon et al. (2017) expands the network using group sparse regularization to decide how many neurons to add at each layer and perform selective retraining. Mallya and Lazebnik (2018) propose to prune the weights with low magnitude and reuse them for the next task leaving the rest that are responsible for the previous tasks unchanged. In architecture-based approaches, typically the old knowledge is not exploited to learn the new, thus preventing knowledge transfer from a related task which is an important factor in continual learning.
Bayasi et al. (2021) proposes a fixed-size architecture-based cl strategy where a subset of the network units is assigned to learn each domain separately. Their approach introduces a novel pruning criterion that allows a fixed network to learn new data domains sequentially over time. They identify culprit units associated with wrong classifications in each domain and use them for learning the new domain while freezing the non-culprit nodes.
González et al. (2022) investigate the performance of continual learning methods in a task-agnostic setting, which better simulates dynamic clinical environments characterized by gradual population shifts. They propose an out-of-distribution detection-based solution that signals when to expand the model and select the best parameters during inference. Specifically, they learn a multivariate Gaussian on the last batch normalization layer of the deep model and store it in the memory. When there is out-of-distribution detection or domain shift is alerted based on a threshold on Mahalnobis distance from existing Gaussian in the memory and the new data, a new Gaussian is added to the memory. The closest Gaussian is used for inference and thus eliminates the need for domain ID during inference.
Karani et al. (2018) propose to use domain/task-specific batch norm layers to adapt to new MRI protocols while learning the segmentation of various brain regions. Entire batch-normalization layers were dedicated to modeling domain differences. However, this causes quadratic parameter increase with new tasks. As this approach dedicates specific batch-normalization parameters to each dataset/domain/task, task labels are necessary that determine to which dataset each sample belongs.
Table 10: List of various hybrid CL technique based works in literature
Reference (year) CL scenario & application CL technique
Ozdemir et al. (2018)
MRI-based Humerus and scapula segmentation: CIS (2 Ep.)
Knowledge distillation + experience replay + multi-head
Ozdemir and Goksel (2019)
MRI based knee segmentation (2 Epi.) CIS
Knowledge distillation + experience replay + multi-head
Ravishankar et al. (2019)
Chest x-ray view classification: DIS (2 Ep.), TIS (2 epi) Pneumothorax identification from X-ray: data-incremental (4 batches, each of 2K)
Latent replay where samples are selected either randomly or importance based (farther from cluster centroids) + dynamic network (task specific feature transformer layers)
Lenga et al. (2020)
Chest x-ray classification: DIS (2 epi)
Comparative study: joint training, EWC, LwF
Li et al. (2020)
Dermoscopic images based skin disease classification: CIS (4-20 Ep.)
Dual knowledge distillation (from old classifier (previous classes) and the fine-tuned classifier (a new FC layer fine-tuned on new classes)) + experience replay (fixed subset of data for each class is stored), during inference nearest-mean-of-exemplars method used for class prediction (Rebuffi et al. 2017) (where the mean is often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class)
Morgado et al. (2021)
DIS: dermatological imaging modality classification
Comparative study: Regularization (EWC), rehearsal based (AGEM, experience replay)
Derakhshani et al. (2022)
TIS, CIS, DIS (4 epi)
Regularization, memory-replay (comparative: EWC, MAS, LwF, iCaRL, EEIL)
Ranem et al. (2022)
MRI based binary hippocampus segmentation: DIS (3 epi)
Comparative study: transformers with EWC fisher matrix (Kirkpatrick et al. 2017), EWC with Riemannian walk (Chaudhry et al. 2018b), Modeling the Background (knowledge distillation and a modified Cross Entropy Loss) (Cermelli et al. 2020), Pseudo-labeling and Local Pod (multi-scale spatial distillation loss with pseudo labeling) (Douillard et al. 2021), Pooled Outputs Distillation (Douillard et al. 2020)
Liu et al. (2022)
CIS: CT data based segmentation of abdomen organs (4 Ep. to sequentially include organs)
Rehearsal (latent replay) + regularization (knowledge distillation)
Akundi and Sivaswamy (2022)
Chest x-ray classification in CIS (5 epi)
Regularization (knowledge distillation) + architecture (task specific head)
Shu et al. (2022)
DIS for fundus disease classification (2-4 epi)
Replay (memory buffer optimization) + regularization (gradient regularization)
Kaustaban et al. (2022)
DIS, CIS, TIS, Data-IL
Regularization, replay (comparative study: EWC, LwF, CoPE, iCaRL, A-GEM)
Chee et al. (2023)
CIS in classification setting for three applications: cancer classification, diabetic retinopathy classification, skin lesions classification
Architecture (separate high-level descriptor for each task) + regularization (distillation loss) + replay (data based rehearsal, fixed memory per task)
Zhang et al. (2023b)
(a) CIS for multi-organ segmentation, (b) CIS for abdomen segmentation to liver tumor segmentation
Architecture: organ specific segmentation head + pseudo labelling based knowledge distillation
Ji et al. (2023)
CIS for multi-organ segmentation
Architecture: organ specific decoder + pseudo labelling based knowledge distillation
Zhang et al. (2023c)
CIS: continual disease classification in different image modality settings
Experience replay + Architecture based (keeps task specific adapters and classifier heads)
Bai et al. (2023)
CIS
Regularization: knowledge distillation
Sadafi et al. (2023)
DIS, CIS, DIS+CIS
Experience replay based rehearsal (representativeness of examples for each class and model uncertainty) + regularization (binary-cross entropy based distillation loss)
Yang et al. (2023b)
MedMNIST based disease classification: CIS+DIS (4 epi)
Generative replay+ data-focused regularization
Bándi et al. (2023)
DIS (cancer detection across organs such as breast, colon, and head-neck)
Comparative study of existing approaches: PackNet (architecture), EWC (regularization), GEM (rehearsal)
4.4 Hybrid techniques
Combining two or more of independent CL strategies has recently gained interest owing to the enhanced performance gained by harnessing the merits of different CL strategies. Ozdemir et al. (2018) mitigate catastrophic forgetting for incremental medical segmentation applications using a distillation loss inspired by LwF Li and Hoiem (2017). Further, they extend the approach by incorporating memory-replay strategies and task-specific segmentation heads. Specifically, representative images were selected based on abstraction layer response and content distance in the last embedding layer. Liu et al. (2022) also use knowledge distillation and a memory buffer to store the prototypical representation of different organ categories. Akundi and Sivaswamy (2022) also keep class-specific classification head along with knowledge distillation to avoid forgetting for chest x-ray classification application where data arrives in class incremental fashion. A pseudo-rehearsal technique was combined with task-specific dense layers for pneumothorax classification Ravishankar et al. (2019) by Ravishankar et al. (2019); however, it suffers from the linear increase of parameters and unbounded memory to store features for every domain.
Li et al. (2020) introduce dual distillation as well as a fixed memory-based experience replay to continually learn the effective model. Specifically, they maintain an expert classifier which is nothing but the previous model fine-tuned on new classes after replacing the old fully-connected layer with a new one. Then a final classifier is distilled using the old classifier and the expert classifier. A fixed amount of old samples is also stored and replayed with current data while training the updated classifier. Then nearest-mean-of-exemplars (often calculated by averaging the feature vectors of the stored or selected subset of data in the feature space for each class) method is used for category prediction at the test time.
Chee et al. (2023) utilize the dynamic expanding network, regularization, as well as data replay for continually learning new classes. A low-level feature extractor is shared across tasks but a high-level feature extractor is especially learned for each subsequent task. Further, an alternate training procedure so as to learn new classes, i.e., newly added high-level feature extractor (by freezing others) and learning the old classes (freezing the new task feature extractor) is followed. A fixed memory per task is kept to replay samples with current samples.
Sadafi et al. (2023) use experience replay-based rehearsal and binary cross-entropy-based regularization terms for mitigating catastrophic forgetting. There is fixed memory and each class has an equal contribution to it. The sampling strategy for exemplar selection for a specific class is twofold: half of the samples are selected based on their distance from the mean; i.e, the closest samples to the class mean are selected, and another half of the samples for the class are selected based on model’s uncertainty for each sample as given in epistemic uncertainty estimation Mukhoti et al. (2021).
Zhang et al. (2023b) exploit knowledge distillation as pseudo-labeling for old classes along with architecture-based CL strategy. In vanilla Swin UNETR architecture they replace the conventional output layer responsible for segmentation with organ-specific light-weight segmentation heads which thus enable independent predictions for any new or previously learned classes. Thus, there is a single encoder and decoder module and multiple organ-specific heads (few MLP layers) on the output of the encoder. In a similar fashion Ji et al. (2023) keep organ-specific encoder and apply knowledge distillation. They also apply network architecture search-based pruning on decoders to maintain network complexity. But in comparison to Zhang et al. (2023b) strategy the network complexity of maintaining multiple decoders is very high.
Zhang et al. (2023c) propose a rehearsal and architecture-based approach. In order to effectively extract discriminative features from a pre-trained feature extractor for different diseases, a learnable lightweight adapter is added between consecutive convolutional stages at each subsequent episode. Furthermore, each task-specific classification head is also added at each step. Then in each episode model aim is to find optimal parameters in the newly added task-specific adapters and classification head using the training data from the new episode (new disease) and preserving a small subset for each previously learned disease.
4.5 Comparative studies
There have been multiple comparative studies of different kinds of CL approaches on various applications like histopathology-based cancer detection Bándi et al. (2023), histopathology-based tumor detection Kaustaban et al. (2022), chest x-ray classification Lenga et al. (2020), dermatological image modality classification Morgado et al. (2021), etc. Lenga et al. (2020) show a comparative study of joint training and CL techniques such as EWC, and LWF to improve model adaptation to new chest X-ray domains arising from cross-sites while mitigating catastrophic forgetting effects. Upon model evaluation on two datasets from different sites (ChestX-ray14 and MIMIC-CXR) referring to domain shifts, the best results are achieved with joint training, but EWC and LWF offer practical solutions as the previous samples need not be stored with the CL approaches. In the digital pathology research field, Kaustaban et al. (2022) systematically evaluate various CL methods, including regularization-based and rehearsal-based approaches on self-augmented domain shift on the H&E dataset. The authors concluded that though regularization-based methods performed well for DIS and Data-incremental scenarios, only the rehearsal-based method (iCaRL) is effective for CIS which is a more challenging scenario. Further, TIS may be even more challenging for digital pathology as compared to other image domains.
Further, with simulated datasets, i.e., MedMNIST (MNIST-like collections of biomedical images) various benchmark CL methods in rehearsal-based (iCaRL), regularization-based (EWC, MAS, LwF), and bias-correction method (EEIL) were evaluated under different CL scenarios Derakhshani et al. (2022). For the majority of the experiments including CIS, TIS, and cross-DIS, the rehearsal-based approach, i.e., iCaRL shows the most promising results for disease classification. Further authors indicated that the existing CL methods which tend to do well in natural image applications (non-medical) may perform inadequately in medical disease classification applications due to their inherent complexity, such as the spatial locality of diseases. Similarly, in the case of a dermatological imaging modality classification problem under domain-shifted condition, Morgado et al. (2021) showed a comparison of EWC, averaged GEM, and experience replay. They concluded that MobileNetV2 with experience replay performs best among the others. González et al. (2023) also provided open-source implementation of five CL strategies including rehearsal, EWC, LwF, Riemannian Walk (RW), and Modeling the Background (MiB), on top of nnU-Net for various CL based applications like hippocampus, prostate, or cardiac segmentation.
Figure 7: A demonstration for rigid and non-rigid task/domain boundaries for a sequence of four domains in a CL pipeline
5 Level of supervision
Here we provide a discussion on what level of supervision is considered and explored in literature while developing a CL method.
5.1 Task or domain ID
Most of the approaches specifically require information about task/domain ID, i.e., which domain/task the current sample belongs to. This information is then used to select or activate the task-specific branch in the model for classification/segmentation/feature extraction. If task/domain ID is needed then the approach is regarded as (a) task/domain aware else (b) task/domain agnostic. However, a well-designed CL approach should not rely on a task/domain ID to perform prediction Aljundi et al. (2017); Bayasi et al. (2021). Most of the approaches presume a-priori knowledge of domain IDs Derakhshani et al. (2022); Zhang et al. (2023b) while some infer from the data Zhang et al. (2023c); Bayasi et al. (2021); Aljundi et al. (2017). A very few do not require this information Derakhshani et al. (2022); González et al. (2022); Perkonigg et al. (2022).
5.2 Task boundaries
Typically, a shift is regarded as a change of data source, i.e., datasets coming from different sites or acquisition environments. Therefore, there are well-separated domain shifts or task boundaries. Thus, the necessary measures to adapt to the changed domain are triggered when there is a change in task/domain. Most of the works in the literature follow the assumption of well-separated task/domain boundaries.
However, in real-life cases, the change in domain/task could be interleaved or smooth rather than an abrupt transition. Such settings come under the umbrella of ‘blurred boundary continual learning’ Wang et al. (2023a). To better visualize the rigid and non-rigid boundary-based sequence of domains, a demonstration is provided via Figure 7. Recently a few works Hofmanninger et al. (2020a); Srivastava et al. (2021); González et al. (2022); Perkonigg et al. (2022) have explored such settings. For example González et al. (2022) consider two kinds of non-rigid shift in hippocampus segmentation datasets. In the first kind, termed ‘shifting source’, they create slowly shifting distributions for the segmentation task with three datasets (HarP, Dryad, and Decathlon) by interleaving samples from these datasets. In another case termed ‘transformed’, they create shift using TorchIO library Pérez-García et al. (2021) which performs various intensity rescaling and affine transform in the Decathlon dataset to create 5 episodes. Another example of a non-rigid boundaries-based CL scenario is given by Srivastava et al. (2021) for the chest x-ray classification problem. They consider three multi-site datasets (NIH Chest-X-rays14, PadChest, and CheXpert) as three domains; however, they curate a smooth transition from one dataset to another at the boundaries.
5.3 Sample annotation
Apart from information regarding the current task, annotation for the samples in the current task is also a major concern. Almost all the available CL strategies in the medical domain require sample annotation for learning the classification or segmentation task. It is undeniable that labeling is highly costly in the medical domain, especially in the histopathology field, and hence may hinder the rapid advancement of the field. In other domains, unsupervised CL strategies have been explored for example Ashfahani and Pratama (2022) only use a few labeled samples to associate clusters to classes and the learning is completely unsupervised manner.
Figure 8: Train-test performance matrix
Table 11: Metrics for backward transfer. For ease of readability, terms in equations are color-coded to refer to the corresponding cells in the matrix (last column).
Reference Eq. Equation Pictorial representation
[…]
Table 12: Metrics for forward transfer. For ease of readability, terms in equations are color-coded to refer to the corresponding cells in the matrix (last column).
Reference Eq. Equation Pictorial representation
[…]
González et al. (2022) Eq.(G) $\frac{1}{T-1}\sum_{i=2}^{T}\left(\frac{1}{\left|\left\{t_{j}\right\}_{j\leq i}\right|}\sum_{j\leq i}\left(P_{j,i}-P_{j-1,i}\right)\right)$
[…]
6 Evaluation Strategy and Metrics
To evaluate a CL strategy various aspects like the performance of the model on past, current, or future episodes, resource consumption Zhang et al. (2023b), memory size Hofmanninger et al. (2020a), model size growth, execution time, etc can be explored. CL frameworks involving dynamic architecture or memory also report time and memory analysis. For example González et al. (2022) maintains model history as mean and variance parameters of multivariate Gaussian which also dynamically grow in the presence of drift; hence authors also provided training time analysis. If the framework does not involve dynamic architecture, mainly performance on data is reported. Further, the performance is mainly reported for evaluating episode performance, stability, and plasticity.
In CL, there is a given sequence of episodes ($1,...,T$) to sequentially learn. Typically we have well-defined training and test pairs as $(Tr_{i},Te_{i})$ for each $i^{th}$ episode. Upon sequential training, a train-test performance matrix $P\in\mathit{P}^{T\times T}$ is generated for the given sequence of episodes ($1,...,T$). An example matrix for $T=4$ is shown in Figure 8. The cell $P_{t,i}$ is performance on test data of $i^{th}$ episode when the training up-to training data of $t^{th}$ episodes is complete. Thus, after training up to $t^{th}$ episodes, the performance $P_{t,i}$ for $t<i$, $t=i$, and $t>i$ correspond to past, current, and future performance, respectively.
The performance metric is chosen depending on the application. For example, IoU, dice score/coefficient, SSIM, Hausdorff distance, and average symmetric surface distance, etc are popular choices for segmentation whereas accuracy, recall, F1-score, and AUC are usually computed for classification applications. Once we have chosen our performance metric, we can compute metrics for measuring stability and plasticity on top of it. Various metrics have been proposed to quantify stability (Table 11) and plasticity (Table 12); however all the metrics are derived from the same train-test performance matrix (Figure 8) to measure the amount of forgetting or forward transfer.
6.1 Stability: Backward Transfer
It refers to the catastrophic forgetting of a model upon learning new episodes. A model is regarded as stable if it gives a similar performance for a particular episode. However, upon learning the new episode naively, it is possible to forget the past learning and hence disrupt the performance of the previous episode. Whereas a carefully designed CL strategy will cause less forgetting or improvement on past episodes upon learning the new episodes. A backward transfer (BWT) metric is used to measure the amount of forgetting. In other words, we can say it is a way to measure stability, i.e., how well the model would retain the previously acquired knowledge to prevent catastrophic forgetting.
Lopez-Paz and Ranzato (2017) define BWT as the influence of new learning $t_{i}$ task on a previous task. When learning a task $t_{i}$ increases the performance on a future task, it is $+ve$ BWT and if it causes deterioration, it results in $-ve$ BWT. Large $-ve$ BWT is also termed as catastrophic forgetting. The BWT metric given by Lopez-Paz and Ranzato (2017) is very frequently adopted Srivastava et al. (2021); Hofmanninger et al. (2020a) and it is computed after finishing all the episodes as shown in Eq.(A) in Table 11. Average forgetting Derakhshani et al. (2021) is another popularly used metric Derakhshani et al. (2022); Chen and Tang (2022). It is computed as the average of the difference in highest performance and the final performance reached after training on all episodes is finished (Eq.(B) in Table 11). Chee et al. (2023) compute BWT after learning each subsequent task ($t^{th}$) rather than computing it at the last episode (Eq.(B) in Table 11) as shown through Eq.(C) in Table 11. González et al. (2022) compute BWT as the change in performance after training with each subsequent task averaged over the number of tasks (Eq.(D) in Table 11). Note that BWT is not defined for the last task. Another BWT metric contribution is by Özgün et al. (2020) as shown via Eq.(E) in Table 11.
6.2 Plasticity: Forward Transfer
The plasticity of a model is reflected in its capability to accommodate more and more knowledge or exploit the knowledge learnt so far to learn the new. In literature, this is frequently referred to as forward transfer (FWT). It is also said that FWT measures the “zero-shot" learning capability of a model Lopez-Paz and Ranzato (2017). In contrast to the usage of BWT, not all works report FWT. Certain approaches might enforce rigorous regularization on the model, which can result in poor performance on new tasks, consequently causing a comparatively reduced level of forgetting, i.e., high $+ve$ BWT. Therefore reporting FWT along with BWT is equally important for unbiased evaluation of the model.
The very popular FWT metric Srivastava et al. (2021); Hofmanninger et al. (2020a) shown via Eq.(F) in Table 12 is proposed by Lopez-Paz and Ranzato (2017) that quantifies FWT as the impact of learning a task $t_{i}$ on a future task. It is the average of the difference in the performance on $t_{i}$ task before learning it and the performance given by a model with random weights ($\bar{b}_{i}$). Note that FWT is not defined for the first task. Further, González et al. (2022) computes FWT as the change in performance in each stage before and upto $Tr_{i}$, averaged over the number of tasks (Eq.(G) in Table 12). Note that it is not defined for the first task. Another metric for FWT was proposed by Shu et al. (2022) as shown via Eq.(H) in Table 12. Özgün et al. (2020) defines two separate forward transfer metrics one for unseen (average of elements above diagonals of $P$) and seen episodes shown in Eq.(I) and Eq.(J) in Table 12, respectively. $b_{i}$ is the performance on a standalone model only trained on $i^{th}$ episode. Further, they report a “transfer-learning metric" as the sum of diagonal elements in matrix $P$ as measures of plasticity, i.e., the ability to adapt to new tasks.
6.3 Episode performance and others
Average accuracy Derakhshani et al. (2021) is computed as the model performance after training up to $t^{th}$ episode. At any time $t$, the average performance is the mean of accuracy values for episodes ($1,...,t$) as below:
$\frac{1}{t}\sum_{i=1}^{t}P_{t,i}$ (1)
Thus, after processing all the episodes, we can get total $T$ average accuracy values which we can plot and visualize for different approaches Derakhshani et al. (2022); Kaustaban et al. (2022). Further, researchers Derakhshani et al. (2022) also compute mean over these values ($\frac{1}{T}\sum_{i=1}^{T}[\frac{1}{t}\sum_{i=1}^{t}P_{t,i}$]) to get a single scale value which can be used to get a single scalar value to directly compare with other approaches. Another popular definition of average accuracy is to compute the above-mentioned metric only after finishing all the episodes Lopez-Paz and Ranzato (2017); Chen and Tang (2022); Kaustaban et al. (2022) as below.
$\frac{1}{T}\sum_{i=1}^{T}P_{T,i}$ (2)
Another performance metric namely “incremental-learning metric" is contributed which is the average of elements in lower-triangular matrix $P$zgün et al. (2020); Li et al. (2022).
Table 13: Literature for CL based classification
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation matrices Dataset
Ravishankar et al. (2019)
Ultrasound, X-ray
Cardiac view, Pneumothorax
DIS, TIS, data-incremental
Rehearsal + architectural
CNN
Supervised
Accuracy, BWT
Subset of ChestXRay (Wang et al. 2017)
Hofmanninger et al. (2020a)
CT
Chest
DIS
Rehearsal
Res-Net50 (He et al. 2016)
Supervised
Accuracy, BWT, FWT
Private
Lenga et al. (2020)
X-ray
Chest
DIS
Comparitive: EWC, LWF
DenseNet121 (Huang et al. 2017)
Supervised
AUC, BWT, FWT
ChestX-ray14, MIMIC-CXR
Li et al. (2020)
Dermoscopic
Skin disease
CIS
Dual-distillation+rehearsal
ResNet18, AlexNet, VGG19
Supervised
Accuracy
Skin8, Skin40 from ISIC2019 (Tschandl et al. 2018), CIFAR100
Srivastava et al. (2021)
X-ray
Chest
DIS
Rehearsal
ResNet101
Supervised
AUC, FWT, BWT
NIH Chest-X-rays14 (McDermott et al. 2020), PadChest (Bustos et al. 2020), CheXpert (Wang et al. 2017)
Bayasi et al. (2021)
Dermoscopic
Skin lesion
DIS
Architectural (fixed but partitioned network)
ResNet-50
Supervised
Accuracy
HAM10000 (Tschandl et al. 2018), Dermofit (Ballerini et al. 2013), Derm7pt (Kawahara et al. 2018), MSK (Codella et al. 2018), PH2 (Mendonça et al. 2013), UDA (Codella et al. 2018)
Morgado et al. (2021)
Dermoscopic
Dermatological imaging modality
DIS
Regularization, rehearsal
VGG-16, MobileNetV2
Supervised
Accuracy, precision, recall, F1-score, FWT, BWT
Private
Shu et al. (2022)
Fundus images
Eyes
DIS
Rehearsal + regularization
Resnet-18, 2-layer MLP (Saha et al. 2021)
Supervised
F1-score, BWT, FWT
ODIR, RIADD, REFUGE, iSee (Fang et al. 2020), EyeQ (Fu et al. 2019), PMNIST (Rupesh Kumar et al. 2013)
Akundi and Sivaswamy (2022)
X-ray
Chest
CIS
Regularization
DenseNet121
Supervised
AUC, BWT
CheXpert
Kaustaban et al. (2022)
Histopathology
Tumor (breast cancer, colorectal cancer)
Data incremental, DIS, CIS, TIS
Regularization, rehearsal (comparative :EWC, LwF, AGEM, CoPE, iCaRL)
ResNet-18
Supervised
Accuracy, FWT, BWT
CRC (Kather et al. 2019), PatchCam (Bejnordi et al. 2017)
Chen and Tang (2022)
Histopathology
Breast cancer
CIS
Regularization
Alexnet, Densenet201, Resnet152
Supervised
Accuracy, forgetting
BreakHis
Derakhshani et al. (2022)
Multiple
Multiple disease
TIS, CIS, cross-DIS
Regularization, rehearsal (comparative: EWC, MAS, LwF, iCaRL, EEIL)
ResNet18
Supervised
Accuracy, forgetting
MedMNIST (Yang et al. 2023a)
Chee et al. (2023)
Histopathology, retinal, skin lesions
Colorectal cancer, eye, skin
CIS
Rehearsal + architecture + regularization
ResNet18
Supervised
Accuracy, forgetting
CCH5000 (Kather et al. 2016), EyePACS (kag ), HAM10000 (Tschandl et al. 2018)
Zhang et al. (2023c)
multiple
Disease (multiple organ source)
CIS
Rehearsal + architectural
ResNet18
Supervised
MCR
Skin8, Path16, CIFAR100
Bándi et al. (2023)
Histopathology
Cancer (multi-organ: breast, colon, head-neck)
DIS
Comparative: (Architecture: PackNet) (Regularization: EWC, GEM)
DenseNet
Supervised
FROC, ROC, Cohen’s Kappa
CAMELYON16, CAMELYON17, private for colon and head-neck
Bai et al. (2023)
VQA
Surgical visual-question localized-answering
CIS
Regularization
VisualBERT
Supervised
Accuracy, IoU
EndoVis18, EndoVis17, M2CAI
Sadafi et al. (2023)
microscopic
WBC
DIS, CIS, DIS + CIS
Rehearsal + regularization
ResNeXt-50
Supervised
Accuracy
Matek-19 (Matek et al. 2019), INT-20, and Acevedo-20 (Acevedo et al. 2020)
Yang et al. (2023b)
Multiple
Disease classification
DIS+CIS
Rehearsal
ResNet18
Few-shot
Accuracy, performance dropping rate
MedMNIST (Yang et al. 2023a) (PathMNIST, DermaMNIST, OrganAMNIST, RetinaMNIST, BreastMNIST, BloodMNIST)
Table 14: Literature for CL based segmentation
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation matrices Dataset
Karani et al. (2018)
MRI
Brain
DIS
Architecture
U-Net
Supervised
DSC
HCP (Van Essen et al. 2013), ADNI (ADNT ), ABIDE (Di Martino et al. 2014) and IXI (IXI )
Baweja et al. (2018)
MRI
Normal brain structures, white matter lesions
TIS
Regularization: EWC
DeepMedic 3D convnet (Kamnitsas et al. 2017)
Supervised
DSC coefficient
UK Biobank (Miller et al. 2016)
McClure et al. (2018)
SMRI
Axial and sagittal brain
DIS
Regularization: DWC
MeshNet (Fedorov et al. 2017) (CNN-based)
Supervised
DSC
HCP (Van Essen et al. 2013), NKI (Nooner et al. 2012), Buckner (Biswal et al. 2010), WU120 (Power et al. 2017)
Ozdemir et al. (2018)
MRI
Humerus and scapula
CIS
Regularization (LwF) + rehearsal
U-Net
Supervised
DSC, SurfDist
Private
Ozdemir and Goksel (2019)
MRI
Knee
CIS
Regularization
U-Net
Supervised
DSC, MSD
SKI10 MICCAI Grand Challenge (Heimann et al. 2010)
van Garderen et al. (2019)
4 MR sequences: pre- and post-contrast T1w, T2w, & T2w FLAIR
Brain tumor
DIS
Regularization: EWC
3D U-Net CNN
Supervised
Dice score
BraTS 2018 (Bakas et al. 2017; Menze et al. 2014), private
Özgün et al. (2020)
MRI
Brain
DIS
Regularization: MAS
QuickNAT (Roy et al. 2019)
Supervised
CL-DSC, REM, BWT+, TL, FWT
CANDI (Kennedy et al. 2012), MALC (Asman and Landman 2013), and ADNI (Jack Jr et al. 2008)
Memmel et al. (2021)
MRI
Hippocampal
DIS
U-Net
Supervised
IoU, DSC
(Simpson et al. ), (Kulaga-Yoskovitz et al. 2015), (Boccardi et al. 2015)
Perkonigg et al. (2021)
MRI
Brain
DIS
Rehearsal
3D-ModelGenesis (Zhou et al. 2021)
Supervised
MAE, FWT, BWT
IXI (IXI ), OASIS3 (LaMontagne et al. 2019)
Ranem et al. (2022)
MRI
Hippocampus
DIS
Comparative (Chaudhry et al. 2018b; Cermelli et al. 2020; Douillard et al. 2021, 2020)
ViT U-Net
Supervised
DSC, FWT, BWT
Decathlon (Antonelli et al. 2022), Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)
Liu et al. (2022)
CT
Liver, spleen, pancreas, right kidney left kidney
CIS
Rehearsal + regularization
nnUNet
Supervised
DC, HD95
(Mul ; Simpson et al. 2019), KiTS (Heller et al. 2019) + private
Karthik et al. (2022)
MRI
Multiple sclerosis lesions (brain)
DIS
Rehearsal
3D U-Net
Supervised
DSC, BWT
(Kerbrat et al. 2020)
González et al. (2022)
MRI
Hippocampus
DIS
Architectural
nnUNet (Isensee et al. 2021)
Supervised
DSC, FWT, BWT
HarP (Boccardi et al. 2015), Dryad (Kulaga-Yoskovitz et al. 2015), Decathlon (Simpson et al. 2019)
Li et al. (2022)
CMR
Cardiac
DIS
Rehearsal
U-Net
Supervised
DSC, HD95, BWT, FWT
M&Ms (Campello et al. 2021)
Zhang et al.; Zhang et al. (2023d; 2021)
MRI
Prostate, optic cup and disc
DIS
Regularization
U-Net
Supervised
DSC, ASD
RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), HCRUDB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (Sivaswamy et al. 2015; Fumero et al. 2011; Orlando et al. 2020)
Zhang et al. (2023b)
CT
Abdomen multi- organ, tumor
CIS
Architectural
Swin UNETR (Hatamizadeh et al. 2021)
Supervised
Average DSC
BTCV (Landman et al. 2015), LiTS (Bilic et al. 2023), JHH (Xia et al. 2022) (private)
Ji et al. (2023)
3D CT
Whole-body organs
CIS
Architectural
nnUNet
Supervised
DSC HD95
TotalSegmentator (Wasserthal et al. 2023), ChestOrgan, HNOrgan, EsoOrgan (3 private)
Bera et al. (2023)
MRI
Prostate, hippocampus, spleen
TIS, DIS
Rehearsal
Residual UNet
Supervised
DSC, average forgetting, BWT, accuracy
(Prostate158 (Adams et al. 2022), NCI-ISBI (NCI ), Promise12 (Litjens et al. 2014), Decathlon (Antonelli et al. 2022)), (Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)), (Spleen dataset in Decathlon (NCI ))
Zhu et al. (2023)
CT, MRI
Abdomen, muscles, prostate
DIS
Regularization
U-Net
Supervised
DSC, ASSD
Prostate (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015), abdomen (Mul ; Kavur et al. 2021), muscles (Zhu et al. 2021)
Liu et al. (2023a)
MRI
Brain tumor
CIS with domain shift
Regularization
ResNet-based 2D nnU-Net
Supervised
DSC, HD95
BraTS2013 (Menze et al. 2014), TCIA (Clark et al. 2013), CBICA (Bakas et al. 2018)
Wang et al. (2023b)
Endoscopic images
Endoscopy, surgical instrument
CIS
Rehearsal + regularization
ResNet101
Supervised
mIoU
EDD2020 (Ali et al. 2021, 2020), EndoVis18 (Allan et al. 2020), EndoVis17 (Allan et al. 2019)
Table 15: CL Literature for classification + segmentation
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation matrices Dataset
Perkonigg et al. (2022)
CMR, CT, MRI
Cardiac segmentation, lung nodule detection, brain age estimation
TIS
Rehearsal
2D-UNet, Faster R-CNN, ResNet-50
Supervised
Dice score, AP, MAE
Cardiac (Campello et al. 2021), LIDC (Setio et al. 2017) + LNDb challenge (Pedrosa et al. 2019), IXI (IXI ) + OASIS-3 (LaMontagne et al. 2019)
7 Discussion: Towards the future
Our exploration of the CL methods for medical imaging reveals a mix of challenges and strategies. The literature reflects a struggle with traditional paradigms, notably the need for explicit task/domain identification. While some methods demand prior knowledge of domain IDs, others explore task/domain-agnostic approaches that aim to eliminate this reliance Aljundi et al. (2017); Bayasi et al. (2021). The evolving notion of task/domain boundaries adds another layer of complexity, with real-world scenarios often defying the assumption of well-separated shifts. Non-rigid transitions and blurred boundaries, as explored by González et al. (2022); Wang et al. (2023a), challenge existing CL approaches. Sample annotation, a common requirement for CL in medical imaging, introduces resource challenges, especially in histopathology. While supervised strategies dominate, Ashfahani and Pratama (2022) showcase the feasibility of unsupervised CL, using minimal labeled samples. The evaluation of CL strategies involves a diverse set of metrics, including stability (backward transfer - BWT) and plasticity (forward transfer - FWT). These metrics, encompassing aspects like average forgetting and incremental learning, facilitate nuanced comparisons. In the classification domain, applications like chest X-rays and skin disease classification employ diverse CL strategies. Latent replay mitigates catastrophic forgetting but increases resource consumption, while knowledge distillation balances stability and plasticity Li et al. (2020); Srivastava et al. (2021). Segmentation tasks, spanning brain sclerosis lesion segmentation to cardiac segmentation, rely on regularization techniques like EWC and knowledge distillation Karthik et al. (2022); Li et al. (2022). The simultaneous handling of classification and segmentation tasks, evident in studies on dermoscopic image-based skin disease classification Li et al. (2020) and abdominal organ segmentation Liu et al. (2022), introduces a trade-off between resource efficiency and overall model performance. The synthesis of literature and comparative analyses unveils the current state of CL in medical imaging. However, challenges persist, and future research should explore unsupervised CL, address labeling costs, and refine evaluation metrics. As the field evolves, understanding the dynamic interplay between medical imaging tasks and CL methodologies will be crucial for driving innovation. in Tables 13, 14 and 15, we present the overview of the selected publications in the domain of classification, segmentation, and both combined, respectively. We highlight the domain/application, dataset information, imaging modalities, continual learning scenarios, and strategies, including the learning process and methodology adopted/proposed by different studies.
Based on our critical analysis of existing literature, we discuss some of the open challenges and thus possible research directions for CL in the medical field.
Availability of benchmark datasets: Contrary to other natural imaging fields, there are no standard benchmark datasets for continual learning in the medical field. Therefore, a fair comparison of different research advancements made in the field is challenging. MedMNIST is an attempt to provide an MNIST-like dataset for the medical field, however, it is limited by its non-complex nature of various sub-datasets and not mainly developed for CL. Hence, it does not offer various drift conditions that a CL technique could be explored to quantify.
Intense labor requirements: Most of the continual learning approaches are supervised in nature and thus demand large annotated datasets for sequential training. This creates a bottleneck in model development especially in the histopathology domain as it is the most expensive in terms of annotation. To cope with a limited labeling budget in problems like cardiac segmentation, lung nodule detection, and brain age estimation, Perkonigg et al. (2022, 2021) explored an active learning strategy within the CL framework. The model indicates important samples and then only those samples are annotated by the experts. On the other hand, there are fully unsupervised CL approaches Ke et al. (2022); Pratama et al. (2021); Ashfahani and Pratama (2022) or self-supervised CL approaches Liu et al. (2023b) in other domains that should be explored here to tackle the annotation scarcity problem.
Unexplored distributional changes: In the medical realm, drifts arising from covariance shifts are mainly explored that too by incorporating datasets from different centers or acquisition protocols for the same task. Thus, the type of drift is always sudden drift and easier to handle. On the other hand, there are many realistic unexplored drift settings. For example, drift within the center is close to a realistic scenario as it gives rise to smooth drift scenarios Perkonigg et al. (2022). Further, some drifts occur temporarily for example, Bichlmayer et al. (2022) observe one novel kind of domain shift in kidney-CAM-model data Bichlmayer et al. (2022), even if they follow a fixed acquisition protocol. The tissue structure in histopathology data of mouse kidneys shows degradation over time which causes the segmentation problem harder over time. Another very important kind of unexplored drift is “concept shift" or “concept drift" which refers to a situation where the shift causes a change in the relationship of the input variable with its target value Lu et al. (2018); Kumari and Saini (2022); Bhatt et al. (2022); Kumari (2021). For example, age group 30-50 are probable to be diabetic patients, but due to unforeseen reasons if the age group shifts to 15-30 then it is a concept drift for the diabetic patient identification task.
Scalability for large-Scale medical datasets: Scalability emerges as a critical challenge in the context of CL for medical imaging, particularly as datasets continue to grow in size and complexity Gonzalez et al. (2020). Efficient strategies are needed to handle the scale of large medical datasets, considering both computational resources and model complexity. Research efforts, as explored by Zhang et al. (2023b), should focus on optimizing model architectures, memory management, and training algorithms to ensure scalability without compromising performance.
Interpretable and explainable models: The interpretability and explainability of CL models constitute another open challenge in the medical imaging domain Rymarczyk et al. (2023). As highlighted in Derakhshani et al. (2022), models used in clinical settings should not only achieve high performance but also provide meaningful insights into decision-making processes. Future research directions should prioritize the development of CL methodologies that can offer explanations for their predictions, fostering trust and acceptance in the clinical community.
Ethical considerations and patient privacy: Ethical considerations and patient privacy concerns are paramount in continual learning for medical imaging. The use of patient data for model training raises ethical questions regarding consent, data anonymization, and potential biases. Future research, in alignment with ethical guidelines discussed by Chee et al. (2023), should prioritize the development of frameworks that ensure responsible data usage, privacy preservation, and ethical guidelines. Striking a balance between innovation and the protection of patient rights is imperative for the sustainable progress of CL in medical imaging
Continual learning in diffusion models: The integration of continual learning into diffusion models Kazerouni et al. (2023) represents a promising avenue for advancing the capabilities of these models in dynamic and evolving environments. Diffusion models, widely employed in diverse fields, including image processing and medical imaging, often face challenges when confronted with changing data distributions. Continual learning techniques can play a pivotal role in allowing these models to adapt and accumulate knowledge over time, ensuring sustained performance in the face of evolving datasets and tasks. By incorporating continual learning, diffusion models can enhance their adaptability and generalizability, making them more resilient to variations in data characteristics and distribution shifts Gao and Liu (2023). This approach can significantly contribute to the robustness and effectiveness of diffusion models in real-world applications.
Continual learning in implicit neural representations: The exploration of continual learning within implicit neural representations Molaei et al. (2023) opens up new frontiers in leveraging the power of neural networks for dynamic learning scenarios. Implicit neural representations, commonly used in generative modeling and function approximation, can benefit immensely from continual learning techniques to adapt to new data and tasks seamlessly. As the data landscape evolves, implicit neural representations face challenges related to retaining past knowledge and efficiently incorporating new information Po et al. (2023); Yan et al. (2021b). Continual learning offers a solution to these challenges by enabling models to update their representations while preserving previously acquired knowledge. This not only enhances the model’s ability to handle changing data distributions but also supports the development of intelligent systems that can learn and evolve over time. Continual learning in implicit neural representations is instrumental for creating adaptable and intelligent models that can navigate the complexities of dynamic environments.
8 Conclusion
It has been proved and accepted by the research community that traditional machine learning models are ill-suited to handle the dynamic nature of data, and continual learning offers a promising solution. The systematic review in this manuscript provides a comprehensive overview of the state-of-the-art research in the field of continual learning in medical imaging analysis. We have explored various aspects of this evolving topic, including the challenges posed by changing data distributions, hardware, imaging protocols, data sources, tasks, and concept shifts in clinical practice and the need for models to adapt seamlessly. Through a meticulous analysis of the existing literature, we have examined the continual learning scenarios, strategies, level of supervision, experimental setup, evaluation schemes, and metrics employed to deal with the drifting nature of data in medical imaging. Furthermore, diverse applications of continual learning in medical imaging are discussed, ranging from disease classification and detection to intricate tasks of image segmentation. Each application area presents unique challenges and opportunities for research and development, and our review has shed light on the progress made in these domains. Additionally, a thorough collection and analysis of various evaluation matrices for forward and backward transfer facilitate robust evaluation and benchmarking of approaches.
In conclusion, this systematic review provides valuable insights into the current state of continual learning adaptation in medical image processing tasks, unexplored challenges, and promising future research directions. As medical imaging technology continues to advance, the development of continual learning models will be crucial for improving diagnostic accuracy, patient care, and overall healthcare outcomes.